ganeshie8 (ganeshie8):

easy proof **using congruences** show that \(\large 2^{4x} \equiv 6 \mod 10 \)

OpenStudy (dan815):

ohh

ganeshie8 (ganeshie8):

for example : \(\large 2^{4*349850934850943850439}\) leaves a remainder \(\large 6\) when divided by \(\large 10\)

OpenStudy (dan815):

2^4x probably has some pattern not the same as showing it must be with mod stuff tho lol

OpenStudy (dan815):

like ending in 16 or

OpenStudy (dan815):

26 36 46...

ganeshie8 (ganeshie8):

exactly! 16*16*16.... always leaves the remainder 6 the question was to prove it using congruences

OpenStudy (dan815):

2 4 8 16<--- 6 32 64 128 256<------ 512 1024 2048 4096<---------- How to show with modularity that every 4th power of 2 must end in 6

OpenStudy (anonymous):

2^(4x) = (2^4)^x = (last digit is 6)^x

OpenStudy (dan815):

ok lol

ganeshie8 (ganeshie8):

a = b mod 10 doesn't mean a^k = b mod 10, right ?

OpenStudy (anonymous):

and last digit of 6^n is always 6

OpenStudy (dan815):

that was too simple haha xD

ganeshie8 (ganeshie8):

Oh okay i see...

OpenStudy (ikram002p):

but why 6^n mod 10 always 6 :P

OpenStudy (dan815):

imagine the long multiplication

OpenStudy (ikram002p):

i mean using congrunence *

ganeshie8 (ganeshie8):

you need to ask that question in abstract algebra/analysis course, not in number theory :P

OpenStudy (dan815):

what does it mean to use congruence?

OpenStudy (ikram002p):

no its in number theory :o abstract has nothing to do with it

ganeshie8 (ganeshie8):

using congruences means to use modular arithmetic properties

OpenStudy (ikram002p):

prove that 6^n= 6 mod 10 6=5+1 maybe work ?

ganeshie8 (ganeshie8):

binomial expansion ?

OpenStudy (ikram002p):

yep , then mod it

OpenStudy (ikram002p):

lol what a word "mod it " xD

OpenStudy (anonymous):

Since 6^2 MOD 10 =6 , 6^n MOD 10 = 6

OpenStudy (ikram002p):

i hope the gost understand this as well :3 @sauravshakya mmm not convising

ganeshie8 (ganeshie8):

A^2 = A => A^n = A

OpenStudy (ikram002p):

since when :0

ganeshie8 (ganeshie8):

since i took LA :) think of idempotent matrices

OpenStudy (ikram002p):

3^2 =9 mod 10 3^3=7 mod 10 counterexample

OpenStudy (ikram002p):

xD teach me LA

ganeshie8 (ganeshie8):

if 3^2 = 3 mod something, then definitely 3^n = 3 mod samething.

OpenStudy (ikram002p):

ohh ok got what u mean

OpenStudy (ikram002p):

wondering what dan was typing :3

OpenStudy (ikram002p):

nice one :o

OpenStudy (dan815):

how come we cant just say that the starting playing must always be a 6

OpenStudy (dan815):